OpenStudy (anonymous):

find the antiderivative of (2x+6)^3

OpenStudy (anonymous):

we havent learned how to find antiderivatives of a function raised to a power so im not quite sure where to start

OpenStudy (tkhunny):

Yes you have. What's the derivative of x^4?

OpenStudy (anonymous):

4x^3

OpenStudy (tkhunny):

So, what is the antiderivative of 4x^3?

OpenStudy (anonymous):

x^4

OpenStudy (tkhunny):

Okay, how about the antiderivative of (x+2)^3?

OpenStudy (anonymous):

would it be (x+2)^4?

OpenStudy (tkhunny):

Very Close! \[\frac{d}{dx}(x+2)^{4} = 4(x+2)^{3}\] See that factor of 4 out in front? This makes the requested antiderivative ¼(x+2)^4. Find the derivative of that and convince yourself.

OpenStudy (anonymous):

ooooo i seeee so the antiderivative of (2x+6)^3 is (1/4)(2x+6)^4